OpenStudy (anonymous):

How do I prove the following function? cot2^x/cscx+1 = 1-sinx/sinx

OpenStudy (jdoe0001):

do you have your trig formula sheet handy?

OpenStudy (anonymous):

My what?

OpenStudy (anonymous):

If you mean my identities, yes.

OpenStudy (jdoe0001):

trigonometry formulas cheatsheet :)

OpenStudy (jdoe0001):

so you can check the trigonometric identities so you can factor them out

OpenStudy (anonymous):

Yeah I have my identities with me.

OpenStudy (jdoe0001):

ok, check the pythagorean identities, specifically the one for csc^2

OpenStudy (jdoe0001):

and from there, get what cot^2 is :)

OpenStudy (anonymous):

I can replace cot^2x with 1-csc^2x

OpenStudy (jdoe0001):

mmm, csc^2-1 you meant

OpenStudy (anonymous):

Oh, my bad.

OpenStudy (anonymous):

csc^2x-1/cscx+1 <--- That's it lol

OpenStudy (jdoe0001):

yes now keep in mind that \[csc(x)^2-1= csc(x)^2-1^2\] which is also the same as \[(csc(x)-1)(csc(x)+1)\]

OpenStudy (anonymous):

(cscx-1)(cscx+1) / cscx+1 So that would simplify to just cscx-1 on the left?

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

Now there's 1-sinx/sinx

OpenStudy (jdoe0001):

and that \[\frac{a-b}{c}=\frac{a}{c}-\frac{b}{c}\]

OpenStudy (anonymous):

1/sin x - sin x/sin x ?

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

So that will end up as just 1/sin x ?

OpenStudy (jdoe0001):

well, sinx/sinx isn't "0"

OpenStudy (anonymous):

1/sinx+1?

OpenStudy (jdoe0001):

\[\frac{1}{sin(x)}+1\]

OpenStudy (anonymous):

So that leaves us at: cscx-1 = 1/sinx +1 ?

OpenStudy (jdoe0001):

yes, check your trig sheet for the "cofunction identities" :)

OpenStudy (anonymous):

I have those.

OpenStudy (jdoe0001):

there is your answer :)

OpenStudy (anonymous):

They all deal with pi/2 though. How do I use them here?

OpenStudy (jdoe0001):

ohh shoot, yes, my bad, I meant "reciprocal identities"

OpenStudy (anonymous):

cscx = 1/sinx,

OpenStudy (anonymous):

But the cscx has a -1 in front of it, and the 1/sinx has a +1

OpenStudy (jdoe0001):

hmmm, a (+) on the right?

OpenStudy (anonymous):

1/sinx +1

OpenStudy (jdoe0001):

well it --> cot2^x/cscx+1 = 1-sinx/sinx <-- is what you had originally

OpenStudy (jdoe0001):

so, can't say where the (+) came from

OpenStudy (anonymous):

1-sinx/sinx 1/sinx - sinx/sinx

OpenStudy (anonymous):

Ohhhhhh nevermind it's a -1

OpenStudy (jdoe0001):

:D

OpenStudy (anonymous):

cscx-1 = 1/sinx -1 Is correct so far, right?

OpenStudy (jdoe0001):

yes

OpenStudy (jdoe0001):

and of course from the "reciprocal identities", that'll turn into the same :)

OpenStudy (anonymous):

and cscx = 1/sinx, so that's it?

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

Thank you! :)

OpenStudy (jdoe0001):

yw